OpenStudy (anonymous):

Solve 4^3x=64^2x-6 without using logs

sam (.sam.):

Can use natural log?

sam (.sam.):

Nevermind there's a trick

sam (.sam.):

\[\Huge 2^{2(3x)}=2^{6(2x-6)}\] \[\Huge 2(3x)=6(2x-6) \] Solve

sam (.sam.):

Make the base same and , you can bring down the exponents

OpenStudy (anonymous):

6x= 12x + 36 -6x=36 x= -6?

OpenStudy (anonymous):

Never mind.

sam (.sam.):

very close, 2(3x)=6(2x−6) 6x=12x-36 -6x=-36 x=6